______
Öx² + 5 - x - 2 = 0
          _____
x = 1 + 6Öx - 9

Answer by AnlytcPhil(1806)

 ______
Öx² + 5 - x - 2 = 0

Isolate the radical
 ______
Öx² + 5 = x + 2

Square both sides:

  ______
(Öx² + 5)² = (x + 2)²

x² + 5 = x² + 4x + 4

     1 = 4x

   1/4 = x

It checks.

=====================================
          _____
x = 1 + 6Öx - 9

Isolate the radical term
          _____
x - 1 = 6Öx - 9

Square both sides

              _____
(x - 1)² = (6Öx - 9)²

x² - 2x + 1 = 36(x - 9)

x² - 2x + 1 = 36x - 324

x² - 38x + 325 = 0

(x - 13)(x - 25) = 0

x - 13 = 0, or x = 13

x - 25 = 0, or x = 25

Both solutions check.

Edwin
AnlytcPhil@aol.com